# Diversity of disc viscosities can explain the period ratios of resonant and non-resonant systems of hot super-Earths and mini-Neptunes

### v1

@@
Migration is a key ingredient for the formation of close-in super-Earth and mini-Neptune systems. The migration rate sets in which resonances planets can be trapped, where slower migration rates result in wider resonance configurations compared to higher migration rates. We investigate the influence of different migration rates - set by the disc’s viscosity - on the structure of multi-planet systems via N-body simulations, where planets grow via pebble accretion. Planets in low viscosity environments migrate slower due to partial gap opening compared to planets forming in high viscosity environments. Consequently, systems formed in low viscosity environments tend to have planets trapped in wider resonant configurations (typically 4:3, 3:2, and 2:1 configurations). Simulations of high viscosity discs mostly produce planetary systems in 7:6, 5:4, and 4:3 resonances. After the gas disc dissipates, the damping forces of eccentricity and inclination cease to exist and the systems can undergo instabilities on timescales of a few 10 Myrs, rearranging their configurations and breaking the resonance chains. We show that low viscosity discs naturally account for the configurations of resonant chains as Trappist-1, TOI-178, and Kepler-223, unlike high viscosity simulations which produce relatively more compact chains. After the gas disc dispersal, about 95% of our low viscosity resonant chains became unstable experiencing a phase of giant impacts. Dynamical instabilities in our low viscosity simulations are more violent than those of high viscosity simulations due to the effects of leftover external perturbers (P200 days). About 50% of our final system ended with no planets within 200 days, while all our systems harbor remaining outer planets. We speculate that this process could be qualitatively consistent with the lack of inner planets in a large fraction of the Sun-like stars. Systems produced in low viscosity simulations alone do not match the overall period ratio distribution of observations, but give a better match to the period distributions of chains, which may suggest that systems of super-Earths and mini-Neptunes form in natal discs with a diversity of viscosities.
@@
In contrast to the solar system, a large fraction of exoplanetary systems harbour super-Earth/mini-Neptune planets in orbits interior to that of Mercury (e.g. Mayor et al. 2011; Fressin et al. 2013; Mulders et al. 2018). These systems of close-in planets normally consist of several planets that can be in- or close-to mean motion resonances (MMRs), e.g. Fabrycky et al. (2014); Pichierri et al. (2019). Furthermore, it seems that the majority of these systems behave like “peas-in-a-pod”, where the planets have roughly equal sizes and an equal distribution of period ratios within each system (Millholland et al., 2017; Weiss et al., 2018; Millholland & Winn, 2021). However, their formation pathway is still under debate.
@@
While it is clear that the formation of systems in resonance requires planetary migration (e.g. Ogihara et al. 2015; Izidoro et al. 2017, 2021; Hühn et al. 2021), it is unclear if the planets originate from beyond the water-ice line or if the planets formed in the inner regions of the disc. This is - of course - independent if the planetary systems are in resonance or not. For selected systems, the chemical composition of these planets could answer this question (e.g. Benneke et al. 2019; Sun et al. 2024), as the migration behaviour sets the composition of these planets (e.g. Bitsch et al. 2019b; Schoonenberg et al. 2019; Bitsch et al. 2021; Izidoro et al. 2022; Chatziastros et al. 2024). However, even with JWST operating and the future ARIEL mission, we are still far away from a bulk characterization of planets in these systems. In addition, a spectroscopic analysis of the planetary atmospheres requires that the planets are transiting and orbit stars close enough to allow a good enough signal-to-noise ratio. However, for most of the Kepler sample this remains a challenge due to the large distance of the Kepler field from Earth, leaving it unclear if the majority of the close-in planets formed interior or exterior to the water ice line from a detailed compositional point of view.
@@
On the other hand, the analysis of the Kepler sample revealed a valley in the radius distribution of these close-in planets (Fulton et al., 2017; Van Eylen et al., 2018), where peaks at 1.5 and 2.0 Earth radii and a dip at 1.8 Earth radii were found. The interpretation of this radius valley is a dichotomy in planetary properties: While the close-in planets with radii less than 1.5 Earth radii are generally thought to be atmosphereless rocky planets, planets with radii larger than 2.0 Earth radii are thought to be mini-Neptunes with rocky cores and hydrogen/helium envelopes or ice dominated planets (e.g. Owen & Wu 2013; Lopez & Fortney 2014; Jin & Mordasini 2018). This indicates a compositional difference between these types of planets and gives hints to a different origin. The radius valley in itself seems to be universal in the sense that it appears at all stellar masses (Ho et al., 2024).
@@
The radius valley can be explained by photoevaporation of close-in planetary atmospheres (e.g. Owen & Wu 2013; Lopez & Fortney 2014 or also by core-powered mass loss (e.g. Gupta & Schlichting 2019). Alternatively, atmospheres of planets could be lost during collisions with other planets (e.g. Liu et al. 2015; Inamdar & Schlichting 2016). The current data supports all of these scenarios and is in particular consistent with a combined scenario, where all these effects can act at the same time (Izidoro et al., 2022).
@@
As the planets migrate inwards, they will stop at the disc’s inner edge (Masset et al., 2006; Flock et al., 2019), where they can then build-up resonance chains (Izidoro et al., 2017; Ogihara et al., 2018; Izidoro et al., 2021). As the gas disc disappears, the damping forces onto to the planets diminish and the resonance chains can be broken by instabilities, which re-arranges the planetary orbits and increase the eccentricities and inclinations of the close-in planets. These instabilities can either be self-triggered (Izidoro et al., 2017, 2021) or can be triggered due to the presence of outer giant planets (e.g. Bitsch & Izidoro 2023). In addition, the instabilities are affected by the proximity of the planets to each other (e.g. Gladman 1993) as well as by their masses. The re-arrangement of the planets within these systems can naturally explain the observed number of transiting planets as well as the general period distribution of these planets (Izidoro et al., 2021), even though the simulations have trouble explain systems in the 2:1 resonance configuration.
@@
Trapping of planets into mean motion resonances depends on the relative migration speed of the planets (e.g. Huang & Ormel 2023). Faster migrating planets can result in more compact systems than slower migrating planets. The migration velocity in itself, however, is set by the disc’s properties, for example, by the disc’s viscosity. At low viscosity even mini-Neptunes can start to open partial gaps that can reduce their migration speed (e.g. Paardekooper et al. 2011; Kley & Nelson 2012; Bitsch et al. 2013). Consequently systems forming in low viscosity environments could harbour wider resonances with consequences for the long-term evolution of these planetary systems. On the other hand, a lower migration speed might imply that the planets are not trapped in resonance at all at the end of the gas-disc phase, because they migrate too slow to reach the inner edge of the disc (e.g. Brasser et al. 2017).
@@
Past simulations of the breaking the chains model focused on systems without giant planets in high viscosity environments (Izidoro et al., 2017, 2021, 2022; Esteves et al., 2020, 2022, 2023) or on systems with outer giants, but in low viscosity environments (e.g. Bitsch et al. 2020; Bitsch & Izidoro 2023). The low viscosity simulations with giant planets already seem to hint at the fact that systems with a 2:1 resonance configuration could be more common in these environments. Here we want to bridge this gap by analysing simulations within the breaking the chains model at low disc viscosity without outer giant planets.
@@
Our work is structured as follows. In section 2 we give a short overview of our model. We show the results of our simulations in section 3 and compare to the high viscosity simulations of (Izidoro et al., 2021). We then discuss our results in the context of other works in section 4 and summarize our work in section 5.
@@ 2 Methods
To study the dynamics and evolution of planetary systems, N-body simulations are needed. For this we employ FLINTSTONE, an N-body framework based on the MERCURY integrator (Chambers, 1999) that includes planetary growth and migration (Bitsch et al., 2019a; Izidoro et al., 2021). We use here the same code that is described in detail in Izidoro et al. (2021) and Bitsch et al. (2019a), and we just repeat here the main ingredients for our simulations.
@@
We initialize our simulations with 30 planetary embryos that are separated by 0.25 AU starting at 2.75 AU in a low viscosity environment ( for all our simulations). In contrast, the simulations by Izidoro et al. (2021) used a viscosity of , so about a factor of 50 larger. We note that the change of viscosity only influences the migration velocity of planets (see below and appendix A), while the other disc parameters (e.g. gas surface density, temperature) are kept - for simplicity - the same. The gas surface density and temperature thus evolve in time exactly in the same way for all out simulations. This approach, while artificial, allows a clean approach to understand how the viscosity influence the formation of planetary chains. This approach is the same as in our previous works (Bitsch et al., 2020; Bitsch & Izidoro, 2023). The planetary embryos of around Moon mass start in an already evolved protoplanetary disc that is 2 Myr old in the framework of Bitsch et al. (2015a), as in our previous work (Bitsch et al., 2019a, 2020; Bitsch & Izidoro, 2023). The embryos have an initial eccentricity and inclination that is randomly assigned ( and degrees).
@@
At the start of the simulations, the planetary embryos start to grow via pebble accretion (see Johansen & Lambrechts 2017 for a review), where we follow the recipe of Johansen et al. (2015). We use an exponentially decaying pebble flux that corresponds to a total of 350 Earth masses over the 3 Myr lifetime of the gas disc. The pebble sizes are limited by drift (Lambrechts & Johansen, 2014) and we reduce the pebble flux by 50% interior of the water ice line to account for water ice evaporation. Furthermore, we then also reduce the pebbles sizes to mm-size in agreement with chondrules in the solar system, as done in our previous work (Bitsch et al., 2019a; Izidoro et al., 2021). Once the planets have reached the pebble isolation mass (Lambrechts et al., 2014; Bitsch et al., 2018b; Ataiee et al., 2018), they stop growing by pebble accretion.
@@
Once planets have reached the pebble isolation mass, where we ignore for simplicity the dependency on viscosity, they should start to accrete gas (Lambrechts et al., 2014; Bitsch et al., 2015b). The exact efficiency of gas accretion is influenced by the planetary core mass as well as by the envelope opacity (Ikoma et al., 2000; Schulik et al., 2019; Bitsch & Savvidou, 2021; Brouwers et al., 2021; Moldenhauer et al., 2021). In addition, efficient atmospheric accretion can be hindered by recycling flows (Cimerman et al., 2017; Lambrechts & Lega, 2017; Moldenhauer et al., 2021; Wang et al., 2023), which transport material that entered the Hill sphere away from the planet before it can be accreted, which is especially efficient in the inner disc regions. As we are primarily interested in the formation of super-Earth/mini-Neptune systems, we chose to not model the gas accretion process onto the planets as in Izidoro et al. (2021). In principle one can interpret this approach as a limit of very high envelope opacity that prevents efficient contraction of the planetary atmospheres, resulting in the formation of planetary cores with a small hydrogen/helium envelope.
@@
As the planets start to grow, they eventually become big enough to start to migrate efficiently via type-I migration, where we follow the formulae from Paardekooper et al. (2011). Once the planets stars to become even larger, they can open gaps in the protoplanetary disc, reducing their migration velocity, where we follow the approach of Kanagawa et al. (2018), as discussed in appendix B. In addition, we apply forces to mimic the eccentricity and inclination damping onto the growing planets, following our previous approaches (Bitsch et al., 2019a; Izidoro et al., 2021). In principle, the damping of eccentricity and inclination is affected by partial gap opening (Pichierri et al., 2023, 2024), which results in a deviation from the damping formulae from Cresswell & Nelson (2008) that we employ in this work. We nevertheless use the same damping formulae as before to be consistent with our previous works and will investigate the influence of the damping formulae in future works. Planets in low viscosity environments can open partial gaps, resulting in slower migration compared to planets in high viscosity environments (see appendix A).
@@
After 3 Myr our protoplanetary discs dissipates on an exponential time-scale, where we reduce the gas surface density uniformly in the disc with an exponential function on a time-scale of 12 kyr within the last 100 kyr. Once the protoplanetary disc has dissipated we integrate the simulations up to 100 Myr to allow for a large enough time for instabilities to occur. During the instabilities, the planets can also collide. We treat collisions as perfect mergers between the planets, due to the fact that even more detailed collisional treatments that include fragmentation (Leinhardt & Stewart, 2012) give the same results in our simulations (Esteves et al., 2022).
@@
At the end of the simulations, we investigate the structure of the planetary systems and also run synthetic observations of the formed systems. The exact method for this is layed out in Izidoro et al. (2021). These synthetic observations are necessary to effectively account for observational bias and compare our results with observations.
@@
The initial conditions (eccentricity, inclination, planetary mass) are all randomized. In order to get a good statistics, we run 50 simulations with slightly varying initial conditions. We show the resulting planetary systems after 3 (end of the gas-disc phase) and 100 Myr of evolution in appendix C.
3 Planetary systems from low viscosity environments
@@
We discuss in this section the outcome of our results. We first investigate the general outcomes of our simulations and compare these to systems formed in the simulations by Izidoro et al. (2021). We then investigate how synthetic observations of these formed systems can explain the period distribution.

### v2

@@
Migration is a key ingredient in the formation of close-in super-Earth and mini-Neptune systems. The migration rate sets the resonances in which planets can be trapped, where slower migration rates result in wider resonance configurations compared to higher migration rates. We investigate the influence of different migration rates —set by disc viscosity— on the structure of multi-planet systems via N-body simulations, where planets grow via pebble accretion. Planets in low-viscosity environments migrate slower due to partial gap opening compared to planets forming in high-viscosity environments. Consequently, systems formed in low-viscosity environments tend to have planets trapped in wider resonant configurations (typically 4:3, 3:2, and 2:1 configurations). Simulations of high-viscosity discs mostly produce planetary systems in 7:6, 5:4, and 4:3 resonances. After the gas disc dissipates, the damping forces of eccentricity and inclination cease to exist and the systems can undergo instabilities on timescales of a few tens of millions of years, rearranging their configurations and breaking the resonance chains. We show that low-viscosity discs naturally account for the configurations of resonant chains, such as Trappist-1, TOI-178, and Kepler-223, unlike high-viscosity simulations, which produce chains that are more compact. Following dispersal of the gas disc, about 95% of our low-viscosity resonant chains became unstable, experiencing a phase of giant impacts. Dynamical instabilities in our low-viscosity simulations are more violent than those of high-viscosity simulations due to the effects of leftover external perturbers (P200 days). About 50% of our final systems end with no planets within 200 days, while all our systems harbour remaining outer planets. We speculate that this process could be qualitatively consistent with the lack of inner planets in a large fraction of the Sun-like stars. Systems produced in low-viscosity simulations alone do not match the overall period ratio distribution of observations, but give a better match to the period distributions of chains, which may suggest that systems of super-Earths and mini-Neptunes form in natal discs with a diversity of viscosities.
@@
In contrast to the Solar System, a large fraction of exoplanetary systems harbour super-Earth/mini-Neptune planets in orbits interior to that of Mercury (e.g. Mayor et al. 2011; Fressin et al. 2013; Mulders et al. 2018). These systems of close-in planets normally consist of several planets that can be in or close to mean-motion resonances (MMRs) (e.g. Fabrycky et al. 2014; Pichierri et al. 2019). Furthermore, it seems that the majority of these systems behave like ‘peas in a pod’, where the planets have roughly equal sizes and there is an equal distribution of period ratios within each system (Millholland et al., 2017; Weiss et al., 2018; Millholland & Winn, 2021). However, their formation pathway is still under debate.
@@
While it is clear that the formation of systems in resonance requires planetary migration (e.g. Ogihara et al. 2015; Izidoro et al. 2017, 2021; Hühn et al. 2021), it is unclear whether the planets originate from beyond the water-ice line or formed in the inner regions of the disc. This is independent of whether or not the planetary systems are in resonance. For selected systems, the chemical composition of the planets could be used to decipher the location of their formation (e.g. Benneke et al. 2019; Sun et al. 2024), as the migration behaviour sets the composition of these planets (e.g. Bitsch et al. 2019b; Schoonenberg et al. 2019; Bitsch et al. 2021; Izidoro et al. 2022; Chatziastros et al. 2024). However, even with JWST operating and the future ARIEL mission, we are still far away from a bulk characterization of planets in these systems. In addition, a spectroscopic analysis of the planetary atmospheres requires that the planets be transiting, and that they orbit stars close enough to allow a sufficient signal-to-noise ratio. However, for most of the Kepler sample, this remains a challenge due to the large distance of the Kepler field from Earth, leaving it unclear as to whether the majority of the close-in planets formed interior or exterior to the water ice line from a detailed compositional point of view.
@@
On the other hand, the analysis of the Kepler sample revealed a valley in the radius distribution of these close-in planets (Fulton et al., 2017; Van Eylen et al., 2018), where peaks at 1.5 and 2.0 Earth radii and a dip at 1.8 Earth radii were found. The interpretation of this radius valley is that it reflects a dichotomy in planetary properties: while the close-in planets with radii of less than 1.5 Earth radii are generally thought to be atmosphereless rocky planets, planets with radii larger than 2.0 Earth radii are thought to be mini-Neptunes with rocky cores and hydrogen–helium envelopes or ice-dominated planets (e.g. Owen & Wu 2013; Lopez & Fortney 2014; Jin & Mordasini 2018). This indicates a compositional difference between these types of planets and may suggest a different origin. The radius valley in itself seems to be universal in the sense that it appears at all stellar masses (Ho et al., 2024).
@@
The radius valley can be explained by photoevaporation of close-in planetary atmospheres (e.g. Owen & Wu 2013; Lopez & Fortney 2014) or by core-powered mass loss (e.g. Gupta & Schlichting 2019). Alternatively, atmospheres of planets could be lost during collisions with other planets (e.g. Liu et al. 2015; Inamdar & Schlichting 2016). The current data support all of these scenarios and are in particular consistent with a combined scenario, where all these effects can act at the same time (Izidoro et al., 2022).
@@
As the planets migrate inwards, they will stop at the inner edge of the disc (Masset et al., 2006; Flock et al., 2019), where they can then build up resonance chains (Izidoro et al., 2017; Ogihara et al., 2018; Izidoro et al., 2021). As the gas disc disappears, the damping forces acting on the planets diminish and the resonance chains can be broken by instabilities, which rearranges the planetary orbits and increases the eccentricities and inclinations of the close-in planets. These instabilities can either be self-triggered (Izidoro et al., 2017, 2021) or can be triggered due to the presence of outer giant planets (e.g. Bitsch & Izidoro 2023). In addition, the instabilities are affected by the proximity of the planets to each other (e.g. Gladman 1993) as well as by their masses. The rearrangement of the planets within these systems can naturally explain the observed number of transiting planets, as well as the general period distribution of these planets (Izidoro et al., 2021), even though the simulations have trouble explaining systems in the 2:1 resonance configuration.
@@
The trapping of planets into MMRs depends on the relative migration speed of the planets (e.g. Huang & Ormel 2023). Faster migrating planets can result in more compact systems than slower migrating planets. The migration velocity in itself, however, is set by the properties of the disc, such as its viscosity. At low viscosity, even mini-Neptunes can start to open partial gaps that can reduce their migration speed (e.g. Paardekooper et al. 2011; Kley & Nelson 2012; Bitsch et al. 2013). Consequently, systems forming in low-viscosity environments could harbour wider resonances, with consequences for the long-term evolution of these planetary systems. On the other hand, a lower migration speed might imply that the planets are not trapped in resonance at all at the end of the gas-disc phase, because they migrate too slowly to reach the inner edge of the disc (e.g. Brasser et al. 2017).
@@
Past simulations of the ‘breaking the chains’ model focused on systems without giant planets in high-viscosity environments (Izidoro et al., 2017, 2021, 2022; Esteves et al., 2020, 2022, 2023) or on systems with outer giants, but in low-viscosity environments (e.g. Bitsch et al. 2020; Bitsch & Izidoro 2023). The low-viscosity simulations with giant planets already seem to suggest that systems with a 2:1 resonance configuration could be more common in these environments. Here we want to bridge this gap by analysing simulations within the breaking the chains model at low disc viscosity without outer giant planets.
@@
This paper is structured as follows. In Sect. 2 we give a short overview of our model. We show the results of our simulations in Sect. 3 and compare these with the results of the high-viscosity simulations of Izidoro et al. (2021). We then discuss our results in the context of other works in Sect. 4 and summarise our conclusions in Sect. 5.
@@ 2 Methods
N-body simulations are needed to study the dynamics and evolution of planetary systems. To this end, we employ FLINTSTONE, an N-body framework based on the MERCURY integrator (Chambers, 1999) that includes planetary growth and migration (Bitsch et al., 2019a; Izidoro et al., 2021). We use the same code here as that described in detail in Izidoro et al. (2021) and Bitsch et al. (2019a), and only summarise the main ingredients of our simulations.
@@
We initialise our simulations with 30 planetary embryos separated by 0.25 AU starting at 2.75 AU in a low-viscosity environment ( for all our simulations). In contrast, the simulations by Izidoro et al. (2021) used a viscosity of , which is about a factor of 50 more viscous. We note that the change of viscosity only influences the migration velocity of planets (see below and Appendix A), while the other disc parameters (e.g. gas surface density and temperature) are kept the same for simplicity. The gas surface density and temperature thus evolve in time exactly in the same way for all our simulations. This approach, while artificial, allows a clean approach to understand how the viscosity influences the formation of planetary chains. This approach is the same as in our previous works (Bitsch et al., 2020; Bitsch & Izidoro, 2023). The planetary embryos of around Moon mass start in an already evolved protoplanetary disc that is 2 Myr old in the framework of Bitsch et al. (2015a), as in our previous work (Bitsch et al., 2019a, 2020; Bitsch & Izidoro, 2023). The embryos have an initial eccentricity and inclination that is randomly assigned ( and degrees).
@@
At the start of the simulations, the planetary embryos start to grow via pebble accretion (see Johansen & Lambrechts 2017 for a review), where we follow the recipe of Johansen et al. (2015). We use an exponentially decaying pebble flux that corresponds to a total of 350 Earth masses over the 3 Myr lifetime of the gas disc. The pebble sizes are limited by drift (Lambrechts & Johansen, 2014) and we reduce the pebble flux by 50% interior of the water ice line to account for water-ice evaporation. Furthermore, we then also reduce the pebbles sizes to millimetre(mm) size, in agreement with chondrules in the Solar System, as done in our previous work (Bitsch et al., 2019a; Izidoro et al., 2021). Once the planets have reached the pebble isolation mass (Lambrechts et al., 2014; Bitsch et al., 2018b; Ataiee et al., 2018), they stop growing by pebble accretion.
@@
Once planets have reached the pebble isolation mass, where we ignore the dependency on viscosity for simplicity, they should start to accrete gas (Lambrechts et al., 2014; Bitsch et al., 2015b). The exact efficiency of gas accretion is influenced by the planetary core mass as well as by the envelope opacity (Ikoma et al., 2000; Schulik et al., 2019; Bitsch & Savvidou, 2021; Brouwers et al., 2021; Moldenhauer et al., 2021). In addition, efficient atmospheric accretion can be hindered by recycling flows (Cimerman et al., 2017; Lambrechts & Lega, 2017; Moldenhauer et al., 2021; Wang et al., 2023), which transport material that entered the Hill sphere away from the planet before it can be accreted, which is especially efficient in the inner disc regions. As we are primarily interested in the formation of super-Earth/mini-Neptune systems, we chose to not model the gas accretion process onto the planets as in Izidoro et al. (2021). In principle, one can interpret this approach as a limit of very high envelope opacity that prevents efficient contraction of the planetary atmospheres, resulting in the formation of planetary cores with a small hydrogen/helium envelope.
@@
As the planets start to grow, they eventually become big enough to start to migrate efficiently via type-I migration, where we follow the formulae from Paardekooper et al. (2011). When the planets start to become even larger, they can open gaps in the protoplanetary disc, reducing their migration velocity, where we follow the approach of Kanagawa et al. (2018), as discussed in Appendix B. In addition, we apply forces to mimic the eccentricity and inclination damping onto the growing planets, following our previous approaches (Bitsch et al., 2019a; Izidoro et al., 2021). In principle, the damping of eccentricity and inclination is affected by partial gap opening (Pichierri et al., 2023, 2024), which results in a deviation from the damping formulae employed in this work, which are from Cresswell & Nelson (2008). We nevertheless use the same damping formulae as used in previous works in the interest of consistency, and will investigate the influence of the damping formulae in future works. Planets in low-viscosity environments can open partial gaps, resulting in slower migration compared to planets in high-viscosity environments (see Appendix A).
@@
After 3 Myr, our protoplanetary disc dissipates on an exponential timescale, where we reduce the gas surface density uniformly in the disc with an exponential function on a timescale of 12 kyr within the last 100 kyr. Once the protoplanetary disc has dissipated, we integrate the simulations up to 100 Myr to allow sufficient time for instabilities to occur. During the instabilities, the planets can also collide. We treat collisions as perfect mergers between the planets, due to the fact that even more detailed collisional treatments that include fragmentation (Leinhardt & Stewart, 2012) give the same results in our simulations (Esteves et al., 2022).
@@
At the end of the simulations, we investigate the structure of the planetary systems and also run synthetic observations of the formed systems. The exact method for this is described in Izidoro et al. (2021). These synthetic observations are necessary to effectively account for observational bias and compare our results with observations.
@@
The initial conditions (eccentricity, inclination, and planetary mass) are all randomised. In the interest of statistical robustness, we ran 50 simulations with slightly varying initial conditions. We show the resulting planetary systems after 3 (end of the gas-disc phase) and 100 Myr of evolution in Appendix C.
3 Planetary systems from low-viscosity environments
@@
In this section we discuss the outcome of our results. We first investigate the general outcomes of our simulations and compare these to systems formed in the simulations by Izidoro et al. (2021). We then investigate how synthetic observations of these formed systems can explain the period distribution.